# Learn, Check, Test – Security Testing Using Automata Learning and Model Checking

### v1

@@ 2.5 Near Field Communication and Integrated Circuit Access
Near Field Communication (NFC) is a communication standard [31, 7] that defines wireless communication for small, powerless embedded systems with low data rates. It defines messages for data transmission (information or I blocks), signaling (supervisory or S blocks), and acknowledgments (receive-ready or R blocks) and the modalities operate communications. The main applications are RFID-tags and wireless chip card access. For the latter, it uses the structures of ISO/IEC 7816-4 [32]. This standard logically segments circuits into applications (comparable to directories in a file system) that can be accessed via Dedicated Files (DFs) and contains Elementary Files (EFs) as data stores. The standard defines commands to access (SELECT) data and applications, as well as manipulate data (READ, WRITE, UPDATE, etc.). It also defines GETCHALLENGE and AUthenTICATE commands to implement challenge-response-based authentication mechanisms that protect sensitive data. Usually, an authentication procedure yields a session key, that is subsequently used to encrypt the access to the protected parts of the system. Every command induces a response that contains an unencrypted two-bytes status code (even if the data itself is encrypted). This code is either 9000 for positive results or 6XXX for various error codes.
@@ 2.6 Electronically Machine-Readable Travel Documents
Electronically Machine-Readable Travel Documents (eMRTDs) are a logical data structure to store travel document (e.g., passport) data on chips. The International Civil Aviation Organization (ICAO) standardizes this format in their Doc 9303 part 10 standard [33]. It defines two DF areas LDS1 (containing the eMRTD application) and LDS2 (containing travel records, visa records, and additional biometrics). The eMRTD contains the the Common (CM), the Country Verifying Certification Authorities (CVCA), and the Document Security Object (SOD) EFs, as well as EFs for 16 data groups that contain various types of data, like personal data, document data, biometrics (fingerprint, iris). The latter are more sensitive and therefore additionally protected, while the rest is protected by either the older Basic Access Control (BAC) or the newer PACE (Password Authenticated Connection Establishment). BAC generates a key based on some cryptographic operationes with the passport number, expiration date, and the owner’s date of birth. PACE uses a password to encrypt a nonce, which then is the base for a Diffie-Hellman-Merkle key exchange that creates a session key. Besides there are some conditional files outside of both LDS, namely Attributes/Info (ATTR/INFO), Directory (DIR), Card Access (CA), and Card Security (CS). Figure 1 shows an overview of the ICAO eMRTD schema.
Figure 1: Logical Data Structure of Machine Readable Travel Documents from  [34]. Amber is the master file (MF), Cyan are dedicated files (DF), Blue are Elementary Files (EF), and Green are key files. Solid frames means mandatory files, dashed ones optional files. Solid boxes donate the LDS contexts, dashed black boxes requirements, and dashed red boxes necessary authentication.
@@ 2.7 Unified Diagnostic Servcies
Unified Diagnostic Services (UDS) is an ISO standard that specifies services for diagnostic testers to control diagnostic (and further) functions on an on-vehicle Electronic Control Unit (ECU) [8]. Therefore, it defines a variety of services, whereby the concrete elaboration mostly lies in the hands of an implementer. The protocol operates on layer 7 of the OSI reference model. Therefore, it relies on an underlying communication protocol. Although it can use multiple protocols, the most proliferated is the Controller Area Network (CAN) protocol. As a client-server-based protocol it defines requests and responses. The requests contain a service ID, optionally followed by sub-service IDs. Particularly, it defines diagnostic and management services: Diagnostic Session Control (0x10), ECUReset (0x11), SecurityAccess (0x27), CommunicationControl (0x28), TesterPresent (0x3E): a keepalive mechanism,, AccessTimingParameter (0x83), SecuredDataTransmission(0x84), ControlDTCSetting (0x85), ResponseOnEvent (0x86), ResponseControl(0x87). It further defines data transmission services: ReadDataByIdentifier (0x22) , ReadMemoryByAddress (0x23) , ReadScalingDataByIdentifier (0x24) , ReadDataByPeriodicIdentifier (0x2A) , DynamicallyDefineDataIdentifier (0x2C) , WriteDataByIdentifier (0x2E) , WriteMemoryByAddress (0x3D), ClearDiagnosticInformation (0x14), ReadDTCInformation (0x19). Lastly also I/O, routine control, and file transfer functions: InputOutputControlByIdentifier (0x2F) , RoutineControl (0x31), RequestDownload (0x34), RequestUpload (0x35), TransferData (0x36), RequestTransferExit (0x37), RequestFileTransfer (0x38). We will use a set of these services as an input alphabet for learning the UDS state machine (see Section 4.2).
3 Generic Security Requirements
In this section, we take some long established security goals [35, 36, 37] that are also aligned with our case studies’ standards documents (ICAO [38] and UDS [39]) and formulate them as checkable requirements in given-when-then [^3] Using given, when, and then as keywords in italics. format [40]. We use the terms MUST, MUST NOT, SHOULD, SHOULD NOT and MAY as defined by the Internet Engineering Task Force (IETF) [41, 42] [^4] Using the same CAPITALIZED style.. These requirements use attributes, we later (in Section 6.1) use as propositions [^5] Using small capitals style. for defining logic properties.
3.1 Authentication
To establish authenticity as a goal, we must require authentication of a user before accessing protected resources (not every resource must be protected, some may be openly accessible – this depends on the system’s security design). As a requirement , we define it as follows:
3.2 Confidentiality
Requirement R2:Given the system is in an arbitrary state, when an unsecured access operation on a protected resource occurs, then it SHOULD NOT be successful.
3.3 Privilege Levels
It is common practice for more complex systems and protocols to define more than one privilege level according to different roles and according privileges of an actor in order to provide access to more critical resources. This also applies for UDS, where this is implemented by different session and security levels [8]. Likewise, for eMRTDs, sensitve data (particularly biometric data like fingerprints and iris scans [43]) have to be additionally protected via terminal authentication according to European extended access control [33, 44].
Requirement R3:Given different privilege levels, when the level of privilege is not sufficient, then a performed operation MUST NOT be successful.
3.4 Key Validity
Another important attribute of secure connections is key validity and recency [^6] To avoid misunderstandings, for this work we use the term to refer to using the most recent key, i.e., no newer key must exist. Since we do not consider a time component in the automata, we do not use recency in the sense of freshness.. An invalid key in this sense is either a wrong key or an old key (not recent [^7] To avoid replay and similar attacks [45, 46]). Obviously, this is a fundament for the prior requirements, since an authentication or encryption that accepts invalid keys is intrinsically flawed.
Requirement R4:Given an authentication OR secure access operation, when an actor is asked for a key, then providing an invalid key MUST NOT be successful.
4 Building the Mealy Machines using Automata Learning
@@
We use Learnlib (see Section 2.2) to create a setup to learn models of different protocol SULs, particularly NFC-based eMRTDs and UDS communications of automotive ECUs. Learnlib provides a variety of learning algorithms of which we use the TTT algorightm [47] for its comparable high performance [18]. We rely on previous work for the learning parts of eMRTDs [34] and UDS [48], respectively. In this work, we annotated the learned models’ states with labels for propositions and translate them into code to allow for model checking. In this section we provide some details of the learning processes for the two protocols. While the learner is the same, the SUL adapters differ due to different input alphabets and different means to access the respective SUL.
@@ 4.1.1 Input Alphabet
The used input symbols for learning eMRTDs contain effectively selecting the dedicated file for LDS1, selecting elementary files (particularly, CA/CVCA, CM, CS/SOD, ATR, DIR, and data groups DG1-16888Selecting CA and CVACA, as well as CS and SOD, respectively, works with the same inputs. The difference is whether the DF for LDS1 is selected or not (see Figure 1)., read binary and update binary [^9] We also tried write binary, search binary, erase binary, and read record. These were not supported on used SULs (error code 0x6D00).. These symbols are used in a plain and a secure (encrypted with a key obtained in an authentication process) version. Furthermore, we defined additionally modified versions of the secured commands both using a wrong key (actually an all-zero key) and an old key from a previous authentication process [^10] Secret keys for secured commands are obtained by successfully carrying out authentication. If no authentication happened in an input trace before using a secured command, an all-zero key is used. The same applies if an old key should be used but none or only a recent one is available.. For authentication, also a basic authentication (BAC) process is triggered via a dedicated input symbol. Since we lack an implementation of another authentication method (e.g., PACE or Terminal Authentication), BAC is the only authentication input we use.
@@ 4.1.2 Interface Device
We use a Proxmark3 NFC interface devices to access eMRTDs [49]. This devices is capable of handling communication with a broad variety of different NFC dialects, as well as reading from, writing on, flashing and emulating them. Besides, it allows for crafting arbitrary frames, altering any part of the data stream to specific needs.
@@ 4.2.3 SUL Adapter
Analogous to the NFC interface, the learner uses a distinctive SUL class. This class communicates with the device adapter program via a socket interface and passes the input symbols forward. It calls a respective input handler function in a python script. This handler function creates a raw frame (i.e., the raw UDS hex sequence encapsulated in a CAN message) corresponding to the input symbol. It then sends it over the system CAN interface and returns the respective response frame to the learner. This python interface builds on previous work [48] and was expanded with some error handling procedures, CAN-FD support and a variety of additional input symbols.
5 Generating the Rebeca Code
Once the learner provides a model of the SUL (stored in GraphViz format), it can be transformed to code in Rebeca for model checking. This happens in two phases: annotating the model with propositions and subsequently translating into code. The model checker also needs properties to check in the form of linear temporal logic (LTL). This section describes how to obtain these three components.
5.1 Annotating the Mealy Machine with Propositions using CPMs
Automata learning provides us with a SUL’s Mealy machine (see Section 2.2). In a Mealy machine we only have inputs and outputs to distinguish different states and each state does not include any other information. Our goal is to determine the security of a system based on certain properties and it is more intuitive to specify the properties based on the information about the states. We therefore annotate the states with relevant propositions. These propositions should provide the necessary context for sensible security checking.
We define intuitive and usable propositions for our context for each state (e.g., authenticated) and provide a method to annotate the model with these propositions in a correct way. Since the only way of determining the propositions is through the system’s behavior (i.e., input and output), a straightforward approach is to define a context-based proposition map (CPM) that describes which input and output would make a propositions true or false. Therefore, we define gain (for becoming true) and lose (for becoming false) conditions. These condition sets ($C$) define when a state gains ($C_{g}$) and losses ($C_{l}$), respectively, a proposition and constitute the CPM. We define a condition ($c\in C$) as a triplet of a set of propositions $S_{P}\in 2^{P}$, an input set $S_{\Sigma}\in 2^{\Sigma}$, and an output set $S_{\Omega}\in 2^{\Omega}$ ($\langle S_{P},S_{\Sigma},S_{\Omega}\rangle$) and $P$ as set of all propositions in the annotated Mealy machine. Which means that a condition could apply to multiple properties ($p\in P$) to be set by one multiple inputs ($\sigma\in\Sigma$) producing one of multiple outputs ($\omega\in\Omega$). The set $C_{g}$ is applied to all transition labels of the model. The target state of each transition gains $S_{P}$ if the transition’s input and output equals $\sigma\in\S_{\Sigma}$ / $\omega\in S_{\Omega}$ of $c$ according to the Mealy machine’s output function $\lambda$. To automatically apply the propagation of propositions, we create an inheritance map for all propositions. That is, we determine for each transition in the system whether it would propagate the proposition in question. This is the case for any transition which input/output does not match $\sigma/\omega$ for a respective proposition $p$ in the lose set $C_{l}$. With this inheritance map, we iteratively determine the cascading of propositions. For each proposition, we start at the states that possess the respective proposition according to $Cg$ and annotate all target states of transitions in the propagation map for the proposition with $p$. With the updated model, we repeat that step until no new states are to annotate. Eventually, we have a state machine annotated with all propositions according to the defined conditions. As a result, we have labeled states like in a Kripke structure. Since we still maintain the Mealy machine labels of transitions, this resulting annotated Mealy machine can be seen as combined Mealy-Kripke structure. We therefore formally define the annotated Mealy Machine $\mathcal{MK}$ as $\mathcal{MK}=(Q,q_{0},\Sigma,\allowbreak\Omega,\allowbreak,\delta,\lambda,\mathcal{L})$, with $Q$ being the set of states, $q_{0}$ the initial state, $\Sigma$ the input alphabet, $\Omega$ the output alphabet, $\delta$ the transition function and $\lambda$ the output function of the Mealy machine. Finally, $\mathcal{L}$ denotes the labeling function that attributes propositions $p\in P$ to the states. $\mathcal{L}$ corresponds to the used CPM.
Figure 2: We need temporary propositions to specify certain properties where the properties are related to labels on the transitions and we do not have a corresponding proposition in the state for them. Here is a general example for a temporary proposition, stemming from a transition output. Left is the original state ($q$), already labeled with a proposition (AUTHAUTH). On the right side, we introduce an internal state ($\tau$) to gain a checkable proposition for a successful read (READOKREADOK). It has this temporary proposition and inherits all other propositions from $q$.
@@ 5.1 Annotating the Mealy Machine with Propositions using CPMs
5.2 Mapping the Annotated Mealy Machine into Rebeca Code
After annotating the Mealy machine with propositions, a fully automated procedure implemented in Java translates the resulting annotated Mealy machine to Rebeca code (see a formal definition in Section 5.3). In the annotated Mealy machine we see input and outputs of each transition. In the corresponding Rebeca code we have two actors: environment (an external entity giving input and collecting output) and a system (the entity whose behavior is defined by the Annotated Mealy machine). The environment actor contains a request message server and message servers corresponding to each output in the annotated Mealy machine. In the request message server we non-deterministically send messages corresponding to each possible input symbol to the system. For these messages related to each input symbol there is message server in the system actor. Each of these message servers calls back the message server of the environment actor that corresponds to the respective output in the respective state (by call back we mean sending a return message). We keep track of each state using a state variable in the system. Additionally, the system entertains state variables for the propositions. Every time a proposition changes from one state to another (gaining or losing) it is set to true or false, respectively, along with changing the state and ? calling the appropriate output message server on the environment actor. Additionally, the algorithm considers the temporary propositions. If $C_{\tau}$ defines a temporary proposition, a corresponding state variable is set on the environment actor. However, all temporary propositions are reset (i.e., set to false) in the environment actor’s request message server, which is subsequently called by the output message server. This assures that the temporary proposition is present in a state for the model checker, but not maintained further, as it is not part of any state of the Mealy machine. As a summary, the main part of the Mealy machine’s logic is modeled in the system actor, while the environment actor can be seen as an external input giver to the system.
5.3 Formal Definition of a Rebeca Template
We formalize the translation of the annotated Mealy machine by defining a formal Rebeca template by substituting the annotated Mealy components into a Rebeca model definition (see Section 2.4). We define this Rebeca model with two actors: $\mathcal{M}:=r_{sys}||r_{env}$$r_{sys}$ is for a system actor that obtains the annotated Mealy machine’s behavior and $r_{env}$ for an environment actor that serves as an input giver. We define the system as $r_{sys}:=\langle V_{sys},M_{sys},K_{sys}\rangle$. We further define $V_{sys}:=\{q\}\cup AP$. So $r_{sys}$ holds a state variable with an identifier of the current state ($q\in Q$ from the annotated Mealy machine, initially set to $q_{0}$) and $P$, which is the annotated Mealy machine’s set of atomic propositions, for each of which $V_{:}{sys}$ contains a boolean state variable. The only known actor is the environment ($K_{sys}:=\{r_{env}\}$) and the method identifiers are equal the set of input symbols ($M_{sys}:=\Sigma$). The environment ($r_{env}:=\langle V_{env},M_{env},K_{env}\rangle$) contains the temporary propositions from the CPM as state variables ($V_{env}:=C_{\tau})$) and system as other known actor ($K_{env}:=\{r_{sys}\}$). The set of method identifiers consists the output alphabet and a request method that handles the calls of the system and resets the temporary propositions ($M_{env}:=\Omega\cup\{req\}$). In each of the methods of $r_{sys}$ ($i\in M_{sys}$) two things happen: a) we set the curent state identifier ($q$) along with all properties ($p\in AP$) according to the respective transition in the annotated Mealy machine’s transition function that corresponds to the current state and input symbol $\delta(q,\sigma)$ [^11] Note: the Method $m\in M_{sys}$ has a direct, bijective representation in the set of Input Symbols ($\sigma\in\Sigma$): $M_{sys}\leftrightarrow\Sigma$. Additionally having the state variable, we can therefore directly use the transition ($\delta$) and output ($\lambda$) functions of the annotated Mealy machine.. b) we call an output method in $r_{env}$ ($oinM_{env}$) according to the current sate and the annotated Mealy machine’s output function $\lambda(q,\sigma)$, which means sending the message $\langle r_{sys},r_{env},M_{env}(\omega)\rangle$. The environment actor’s ($r_{env}$) request method (reqreq), resets all state variables (which represent the temporary propositions from $C_{\tau}$) to false and randomly picks one input by calling a method from $M_{sys}$ by sending the message $\langle r_{env},r_{sys},M_{sys}(\sigma)\rangle$, with a randomly picked $\sigma\in\Sigma$. Each of the other methods (mapping to a $\omega$ from the annotated Mealy machine) sets a state variable corresponding to a temporary proposition to true, if there is a respective entry in the CPM ($C_{\tau}$) [^12] Depening on the calling methods, which is submitted as a parameter in the message. and then calls the request method (with the message $\langle r_{env},r_{env},{req}\rangle$) to start another interaction cycle. Lastly, we note that the request method is also called by $r_{env}$’constructor, which starts the execution of the Rebeca model ($\mathcal{M}$.
5.4 Illustrative Example

### v2

@@ 2.5 Near Field Communication and Integrated Circuit Access
It also defines GETCHALLENGE and AUthenTICATE commands to implement challenge-response-based authentication mechanisms that protect sensitive data. Usually, an authentication procedure yields a session key, that is subsequently used to encrypt the access to the protected parts of the system. Every command induces a response that contains an unencrypted two-bytes status code (even if the data itself is encrypted). This code is either 9000 for positive results or 6XXX for various error codes.
@@ 2.6 Electronically Machine-Readable Travel Documents
Electronically Machine-Readable Travel Documents (eMRTDs) are a logical data structure to store travel document (e.g., passport) data on chips. The International Civil Aviation Organization (ICAO) standardizes this format in their Doc 9303 part 10 standard [16]. It defines two Dedicated File (DF) areas LDS1 (containing the eMRTD application) and LDS2 (containing travel records, visa records, and additional biometrics). The eMRTD contains the Common (CM), the Country Verifying Certification Authorities (CVCA), and the Document Security Object (SOD) Elementary Files (EFs), as well as EFs for 16 data groups that contain various types of data, like personal data, document data, and biometrics (fingerprint, iris). The latter are more sensitive and therefore additionally protected, while the rest is protected by either the older Basic Access Control (BAC) or the newer PACE (Password Authenticated Connection Establishment). BAC generates a key based on some cryptographic operations with the passport number, expiration date, and the owner’s date of birth. PACE uses a password to encrypt a nonce, which then is the base for a Diffie-Hellman-Merkle key exchange that creates a session key. Besides there are some conditional files outside of both LDS, namely Attributes/Info (ATTR/INFO), Directory (DIR), Card Access (CA), and Card Security (CS). Figure 2 shows an overview of the ICAO eMRTD schema.
Figure 2: Logical Data Structure of Machine Readable Travel Documents from  [29]. This shows the applications and files on an eMRTD chip. Amber is the master file (MF), Cyan are dedicated files (DF), Blue are Elementary Files (EF), and Green are key files. Solid frames means mandatory files, dashed ones optional files. Solid boxes donate the LDS contexts, dashed black boxes requirements, and dashed red boxes necessary authentication.
@@ 2.7 Unified Diagnostic Servcies
Unified Diagnostic Services (UDS) is an ISO standard that specifies services for diagnostic testers to control diagnostic (and further) functions on an on-vehicle Electronic Control Unit (ECU) [19]. Therefore, it defines a variety of services, whereby the concrete elaboration mostly lies in the hands of an implementer. The protocol operates on layer 7 of the OSI reference model. Therefore, it relies on an underlying communication protocol. Although it can use multiple protocols, the most proliferated is the Controller Area Network (CAN) protocol. As a client-server-based protocol it defines requests and responses. The requests contain a service ID, optionally followed by sub-service IDs. Table 1 shows an overview of some of these services and their identifiers.
Table 1: Some of the most widely used UDS services.
Service Identifier
Diagnostic Session Control 0x10
ECUReset 0x11
ClearDiagnosticInformation 0x14
ReadDTCInformation 0x19
ReadDataByIdentifier 0x22
ReadMemoryByAddress 0x23
ReadScalingDataByIdentifier 0x24
SecurityAccess 0x27
CommunicationControl 0x28
ReadDataByPeriodicIdentifier 0x2A
DynamicallyDefineDataIdentifier 0x2C
WriteDataByIdentifier 0x2E
InputOutputControlByIdentifier 0x2F
RoutineControl 0x31
RequestDownload 0x34
RequestUpload 0x35
TransferData 0x36
RequestTransferExit 0x37
RequestFileTransfer 0x38
WriteMemoryByAddress 0x3D
TesterPresent 0x3E
AccessTimingParameter 0x83
SecuredDataTransmission 0x84
ControlDTCSetting 0x85
ResponseOnEvent 0x86
ResponseControl 0x87
We will use a set of these services as an input alphabet for learning the UDS state machine.
3 Building Mealy Machines using Automata Learning
@@
We use Learnlib to create a setup to learn models of different protocol SULs, particularly NFC-based eMRTDs and UDS communications of automotive ECUs. Learnlib provides a variety of learning algorithms of which we use the TTT algorightm [20] for its comparable high performance [54]. We rely on previous work for the learning parts of eMRTDs [29] and UDS [10], respectively. In this work, we annotated the learned models’ states with labels for propositions and translate them into code to allow for model checking. As black-box-based automata learning of real-world systems naturally introduces the possibility of errors into the learned model. For that reason reason we treat found violations in the model not as issues but test cases only. We then use this test case on the real system. If the model behaves wrongly as predicted, the violation was confirmed. Otherwise, we gained a counterexample to feed the learner for a refined model. In this section we provide some details of the learning processes for the two protocols. While the learner is the same, the SUL adapters that provide access to the physical system, differ due to different input alphabets and different means to access the respective SUL.
@@ 3.1.1 Input Alphabet
The used input symbols for learning eMRTDs contain effectively selecting the dedicated file for LDS1, selecting elementary files (particularly, CA/CVCA, CM, CS/SOD, ATR, DIR, and data groups DG1-16, read binary and update binary. Selecting CA and CVACA, as well as CS and SOD, respectively, thereby works with the same inputs. The difference is whether the DF for LDS1 is selected or not (see Figure 2). We also tried write binary, search binary, erase binary, and read record. These were not supported on used SULs (error code 0x6D00). These symbols are used in a plain and a secure (encrypted with a key obtained in an authentication process) version. Furthermore, we defined additionally modified versions of the secured commands both using a wrong key (actually an all-zero key) and an old key from a previous authentication process. Secret keys for secured commands are obtained by successfully carrying out authentication. If no authentication happened in an input trace before using a secured command, an all-zero key is used. The same applies if an old key should be used but none or only a recent one is available. For authentication, also a basic authentication (BAC) process is triggered via a dedicated input symbol. Since we lack an implementation of another authentication method (e.g., PACE or Terminal Authentication), BAC is the only authentication input we use.
@@ 3.1.2 Interface Device
We use a Proxmark3 NFC interface devices to access eMRTDs [13]. This device is capable of handling communication with a broad variety of different NFC dialects, as well as reading from, writing on, flashing and emulating them. Besides, it allows for crafting arbitrary frames, altering any part of the data stream to specific needs.
@@ 3.2.3 SUL Adapter
Analogous to the NFC interface, the learner uses a distinctive SUL class. This class communicates with the device adapter program via a socket interface and passes the input symbols forward. It calls a respective input handler function in a Python script. This handler function creates a raw frame (i.e., the raw UDS hex sequence encapsulated in a CAN message) corresponding to the input symbol. It then sends it over the system CAN interface and returns the respective response frame to the learner. This Python interface builds on previous work [10] and was expanded with improved error handling, CAN-FD support and around the double amount of input symbols from the UDS protocol.
Figure 3: A graphical Mealy machine representation of an example for a simple authenticated file read mechanism.
4 Context-Based Proposition Maps
Once the learner provides a model of the SUL (stored in GraphViz format), we create Context-Based Proposition Maps (CPMs) to annotate the model with propositions of our interest. The generic requirements we defined in this paper are the basis for an engineer to populate the CPM table. We defined generic properties (Section 7.2) which constitute of generic propositions like authentication and successful access. The engineer makes these generic propositions specific in the CPM table based on the specific protocol. Each state in the Mealy machine is augmented with propositions which will also be included in the Rebeca model and will be used for checking the desired properties while model checking.
4.1 Motivating Example
To give an illustrative example, we define an Mealy machine that represents a trivial authenticated file access. The machine has two states ($Q=\{S0,S1\}$), two input ($\Sigma=\{$CRED,READ}\}), and two output ($\Omega=\{$OK,ERR}\}) symbols. Further, we have the initial state $q_{0}=S0$, the transition function
@@ 4.1 Motivating Example
and the output function